# A kernel conditional two-sample test

### v1

@@ A.1 Comparing other properties of conditional distributions
We illustrate the preceding ideas by showcasing how the choice of kernel $\kappa$ influences what is being tested, and the power of the test for a fixed amount of data.
@@
We consider the Gaussian kernel $k$ on $\mathcal{X}=[-1,1]^{2}$ with bandwidth $\gamma^{2}=0.25$, pick a function in the RKHS of $k$, and generate measurements with two different additive noises: one is $\mathcal{N}(0,0.025^{2})$, and the other is a mixture of $\mathcal{N}(-\mu,0.025^{2})$ and $\mathcal{N}(\mu,0.025^{2})$, for different values of $\mu$. This choice guarantees that the noise distributions have both zero means, but differ for higher moments. We embed the measurements in $\mathcal{G}=\mathcal{H}_{\kappa}$, where $\kappa$ is either the inhomogeneous linear or the Gaussian kernel with bandwidth $\gamma^{2}=0.05$. The corresponding conditional expectations in $\mathcal{G}$ differ for the Gaussian kernel (since it is characteristic [26]), but coincide for the linear kernel. As a result, a test with such data should trigger with $\kappa$ the Gaussian kernel, and not trigger with $\kappa$ the linear one. We run vector-valued KRR with kernel $K=k\cdot\mathrm{id}_{\mathcal{H}_{\kappa}}$ on data sets with $100$ points, bootstrap test thresholds, and run the test on the locations present in the data set. Figure 3 shows the resulting empirical positive rate for different $\alpha$-levels, confirming that the Gaussian kernel confidently sees a difference between the distributions, whereas the linear kernel only compares means and therefore does not reject more often than by chance. The results are averaged over 100 runs of the experiment, with the shaded regions reporting the 2.5% and 97.5% quantiles.
@@
However, a rich kernel like the Gaussian one can be less efficient at comparing lower moments of distributions. We show this by comparing empirical error rates of the test for inhomogeneous polynomial kernels of different degrees and Gaussian kernels of different bandwidths for $\kappa$. We use the same setup as before, but repeat this for many independently generated data sets. For the type $\mathrm{I}$ error, we use the same function from $\mathcal{X}$ to $\mathcal{Z}$ and noise distribution to generate both data sets. For the type $\mathrm{II}$ error, we use different functions from $\mathcal{X}$ to $\mathcal{Z}$, but the same additive Gaussian noise distribution, so that the distributions on $\mathcal{Z}$ differ by their means but higher moments coincide. Figure 4 shows that rich kernels, seeing fine-grained distributional differences (like the Gaussian one with small bandwidth), are less efficient at recognizing differences in the mean than less rich kernels (e.g., linear kernel) tailored to that particular moment of the distribution.
@@
Figure 3: Comparing the Gaussian and inhomogeneous linear kernels on different distributions having the same conditional mean on the measurement set $\mathcal{Z}$. A trigger is a true positive for the Gaussian kernel, and a false positive for the linear one.
@@
Figure 4: Empirical type $\mathrm{I}$ and $\mathrm{II}$ errors for different output kernels. For the type $\mathrm{II}$ error, conditional expectations in $\mathcal{Z}$ differ but higher moments coincide. We see that Gaussian kernels are more conservative and achieve higher type $\mathrm{II}$ error.
A.2 Summary of algorithms
Our main test procedure is summarized in Algorithm 1, with threshold calibration via bootstrap shown in Algorithm 2. Both routines assume the form $K=k\cdot\mathrm{id}_{\kappa}$ described in the previous section. Then, the test statistic $\lVert f_{D_{1},\lambda_{1}}(x)-f_{D_{2},\lambda_{2}}(x)\rVert_{\mathcal{G}}$ for the test (7) is the norm of the difference between conditional mean embeddings, and is called the conditional maximum mean discrepancy (CMMD) [26]. We provide its computation in Algorithm 3.
Algorithm 1 Kernel conditional two-sample test
1:data sets $D_{1},D_{2}$; kernels $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ and $\kappa:\mathcal{Z}\times\mathcal{Z}\to\mathbb{R}$; tested covariate region $\mathcal{S}\subseteq\mathcal{X}$; bootstrap covariate region $\mathcal{M}\subseteq\mathcal{X}$; significance level $\alpha$, regularization parameter $\lambda$
3:$\tilde{\beta}_{1}\leftarrow\mathrm{Bootstrap}(D_{1},\mathcal{M},\alpha/2,\lambda)$
4:$\tilde{\beta}_{2}\leftarrow\mathrm{Bootstrap}(D_{2},\mathcal{M},\alpha/2,\lambda)$
5:$\textbf{return}{}\{x\in\mathcal{S}\mid\mathrm{CMMD}(D_{1},D_{2},k,\kappa,\lambda,x)>\tilde{\beta}_{1}\cdot\sigma_{D_{1},\lambda}(x)+\tilde{\beta}_{2}\cdot\sigma_{D_{2},\lambda}(x)\}$
Algorithm 2 Bootstrap
1:$D$; $\mathcal{M}$; $\alpha$; $\lambda$
2:Chooses $\tilde{\beta}_{\lambda}$ s.t. test does not reject on $\mathcal{M}$ for an $(1-\alpha)$-fraction of bootstrapped data sets
3:Initialize empty array $\mathcal{B}[1..M]$triangleright $M\gg 1$ is the number of resamples
4:for $j\leftarrow 1$ to $M$ do
5:     $D_{1},D_{2}\sim D$triangleright Resample $D_{1},D_{2}$ from $D$ uniformly with replacement
6:     $\mathcal{B}[j]\leftarrow\max_{x\in\mathcal{M}}\left(\frac{\mathrm{CMMD}(D_{1},D_{2},k,\kappa,\lambda,x)}{\sigma_{D_{1},\lambda}(s)+\sigma_{D_{2},\lambda}(s)}\right)$triangleright Smallest $\tilde{\beta}$ s.t. test does not reject on $D_{1},D_{2}$
7:return $(1-\alpha)$-quantile of $\mathcal{B}$
Algorithm 3 CMMD: Conditional maximum mean discrepancy
1:data sets $D_{1},D_{2}$; covariate $x\in\mathcal{X}$, $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$, $\kappa:\mathcal{Z}\times\mathcal{Z}\to\mathbb{R}$, $\lambda$
3:$\mathbf{K}_{i}\leftarrow[k(x_{j},x_{h})]_{1\leq j,h\leq\mathrm{len}(D_{i})},\quad i\in\{1,2\}$
4:$\mathbf{k}_{i}\leftarrow[k(x_{j},x)]_{1\leq j\leq\mathrm{len}(D_{i})}^{\top},\quad i\in\{1,2\}$
5:$\mathbf{a}_{i}\leftarrow\mathrm{solve}_{\mathbf{x}}\left((\mathbf{K}+\lambda I_{\mathrm{len}(D_{i})})\mathbf{x}=\mathbf{k}_{i}\right),\quad i\in\{1,2\}$
6:$\mathbf{C}_{ij}\leftarrow[\kappa(z_{h},z_{r})]_{1\leq h\leq\mathrm{len}(D_{i})}^{1\leq r\leq\mathrm{len}(D_{j})},\quad i,j\in\{1,2\}$
7:return $\left(\mathbf{a}_{1}^{\top}\mathbf{C}_{11}\mathbf{a}_{1}-2\cdot\mathbf{a}_{1}^{\top}\mathbf{C}_{12}\mathbf{a}_{2}+\mathbf{a}_{2}^{\top}\mathbf{C}_{22}\mathbf{a}_{2}\right)^{1/2}$
A.3 Process monitoring
Our test can be used to detect changes in time series, where “change” is understood as the change of the conditional probability of future values conditioned on the past (or on the values of other parameters). In short, the method consists in testing between a fixed reference data set and a sliding window. We illustrate this by detecting changes in the dynamics of a linear system. We emphasize, however, that the purpose of this section is not to claim superiority over other methods for change detection, but rather to showcase how these new tools can be used to achieve this objective. We leave comparison and benchmarking to future work.
We set $\mathcal{X}=\mathbb{R}^{d}$, $d\in{\mathbb{N}_{>0}}$, and choose the linear operator-valued kernel $K(x,x^{\prime})=\langle x,x^{\prime}\rangle_{\mathbb{R}^{d}}\cdot\mathrm{id}_{\mathbb{R}^{d}}$, as we wish to compare dynamics functions mapping $\mathbb{R}^{d}$ to itself. We randomly sample a matrix $A\in\mathrm{O}(d)$, the group of orthonormal matrices in dimension $d$, pick a fixed initial state $X_{0}=\frac{1}{\sqrt{d}}[1\cdots 1]^{\top}\in\mathbb{R}^{d}$, and generate a single trajectory $(X_{n})_{n\in\left[N\right]}$ as $X_{n+1}=AX_{n}+\epsilon_{n}$, where $\epsilon_{n}\sim\mathcal{N}(0,\sigma^{2}I_{d})$ is i.i.d. Gaussian noise with $\sigma=0.01$. We enforce that $A$ is orthonormal so the trajectory neither shrinks nor explodes exponentially. Then, the observations $Y=(Y_{n})_{n\in{\mathbb{N}_{>0}}}$ are taken as $Y_{n}=X_{n+1}$. It is immediate to verify that $D=((X_{n},Y_{n}))_{n\in{\mathbb{N}_{>0}}}$ is a process of transition pairs of the Markov kernel $p(\cdot,x)=\mathcal{N}(\cdot\mid Ax,\sigma^{2}I_{d})$, where $\mathcal{N}(\cdot\mid\mu,\Sigma)$ is the Gaussian measure on $\mathbb{R}^{d}$ with mean $\mu\in\mathbb{R}^{d}$ and covariance matrix $\Sigma\in\mathbb{R}^{d\times d}$.
Using this procedure for $5$ independent system trajectories of length $400$, we collect a data set of $N_{\mathrm{ref}}=2000$ data points, which we call the reference data set. We then follow the trajectory with a sliding window of size $N_{\mathrm{win}}=50$ data points, and perform the test at every step on the points in the sliding window. After $200$ steps, we introduce a change in the matrix $A$, replacing it with the matrix $A^{\prime}=A\exp(\frac{\xi}{\|H\|_{1}}H)$. Here $H=\frac{1}{2}(B-B^{\top})$ is antisymmetric and the entries of $B$ are sampled i.i.d. from $\mathcal{N}(0,1)$, such that $\mathrm{dist}_{\mathrm{O}(d)}(A,A^{\prime})=\xi$, where $\mathrm{dist}_{\mathrm{O}(d)}(A,A^{\prime})=\|\log(A^{\top}A^{\prime})\|_{1}$ is the usual geodesic distance on $\mathrm{O}(d)$. This construction ensures that $A^{\prime}$ is orthonormal as well, and controls the distance between $A$ and $A^{\prime}$ in a way consistent with the geometry of $\mathrm{O}(d)$.
Figure 5 (top) shows the result of this procedure, for varied values of $d$ and of $\xi$ ($\xi$ is only varied with $d=16$), and averaged over $100$ independent choices of the data sets and $50$ independent choices of the matrix $A$. We observe that the change is detected rapidly, even before the sliding window is entirely filled with data from the new dynamics. Notably, higher dimensions require higher disturbances to be successfully detected with a fixed amount of data. Figure 5 (bottom) enables hypothesizing a reason for this: the variance of the reference data set at the current state increases after change, indicating that the dynamics with $A^{\prime}$ bring the trajectory into regions unexplored when following the dynamics with $A$. In other words, the lack of power in high dimensions seems to result from a lack of local data, as it is “easier” to go in unsampled regions by changing the dynamics.
Figure 5: Top: Ratio between the test statistic and the threshold for trigger when monitoring trajectories of a switched linear system in various dimensions and for various distances between the dynamics before and after switch (controlled by $\xi$, shown by the dotted lines). The test triggers when the ratio exceeds $1$. Solid or dotted lines are means; confidence regions show the $5\%$ and $95\%$ quantiles. Bottom: Ratio between the standard deviation $\sigma_{D_{t},\lambda}(x)$ of the sliding window $D_{t}$ and $\sigma_{D_{0},\lambda}(x)$, averaged over $x\in D_{t}$. The variance increases over time, indicating that fewer data points of $D_{0}$ are close to the evaluation points and that the new dynamics are steering the system in unseen regions.
@@ Appendix B Bootstrapping
These considerations strongly support relying on an estimation scheme such as bootstrapping to obtain suitable values for $\beta_{1}$ and $\beta_{2}$ in practice. There are two main difficulties in providing such a scheme: it needs to account for the regions $\mathcal{S}$, $\mathcal{N}_{1}$, and $\mathcal{N}_{2}$ in which we wish to have guarantees, and there are two values to bootstrap. While the first difficulty is easily addressed by sampling (e.g., gridding, random sampling), the second one prevents the boostrapping of a single statistic. In short, we address it by estimating a suitable value for the constant $\beta$ under the null hypothesis for each system separately.
@@
In what follows, we assume that $D_{1}$ and $D_{2}$ are the data sets effectively obtained when sampling the processes of transition pairs of $p_{1}$ and $p_{2}$, respectively. We let $n_{1}=\mathrm{len}(D_{1})$ and $n_{2}=\mathrm{len}(D_{2})$. Furthermore, we assume that we have a sufficiently dense set of points =$\mathcal{M}=\{x_{j}\mid j\in\left[J\right]\}\subset\mathcal{S}$. Such a set can be obtained for instance via gridding or i.i.d. sampling of a distribution on $\mathcal{S}$. We neglect bootstrapping over time, meaning that the values obtained are only meaningful for $(\mathcal{S},n_{1},n_{2})$-guarantees. Finally, we assume that we make the choice $B_{i}=\beta_{i}\cdot\sigma_{D_{n_{i}}^{(i)},\lambda_{i}}$, such that the test $\mathcal{T}$ is effectively parameterized by the vector $\beta=(\beta_{1},\beta_{2})\in\mathbb{R}_{>0}$. Our goal is finding numerically a minimally conservative value for $\beta$ as measured by the empirical type $\mathrm{II}$ error on test scenarios while preserving a type $\mathrm{I}$ error of at most $\alpha$.
Method
@@
For each $i\in\{1,2\}$, we resample with replacement $M$ pairs of data sets from $D_{i}$, where each data set in the pair has the same length as $D_{i}$. Let us call them $((\Delta_{m},\Delta^{\prime}_{m}))_{m\in\left[M\right]}$. For each $m\in\left[M\right]$, then compute $\bar{\beta}_{i,m}$ defined as the minimum $\bar{\beta}_{m}\in\mathbb{R}_{>0}$ such that
@@
$\left\lVert f_{\Delta_{m},\lambda_{1}}^{(1)}(x)-f_{\Delta^{\prime}_{m},\lambda_{2}}^{(2)}(x)\right\rVert\leq\bar{\beta}_{m}(\sigma_{\Delta_{m},\lambda_{1}}(x)+\sigma_{\Delta^{\prime}_{m},\lambda_{2}}(x)),$
for all $x\in\mathcal{M}$. Finally, if $i=1$, define $\beta_{i}\in\mathbb{R}_{>0}$ as the $1-\alpha t$ quantile of the resulting distribution, with $t\in(0,1)$ a user-specified parameter. If $i=2$, pick the $1-\alpha(1-t)$ quantile instead. If it is necessary, a grid search over the parameter $t\in(0,1)$ is cheap. We used $t=\frac{1}{2}$ in all of our simulations.
@@
Figure 6: Influence of the data set size $N$ on bootstrapping performance. The results are averaged over 100 runs of the experiment, with the shaded regions reporting the 2.5% and 97.5% quantiles.
We now investigate whether our bootstrap approach behaves reasonably by investigating the influence of the size of the data set and of the number of bootstrapped data sets on performance.
For this, we set $\mathcal{X}=[-1,1]^{2}$ and $k$ as the Gaussian kernel on $\mathcal{X}$ with bandwidth $\gamma^{2}=0.25$, and generate noisy measurements of functions in $\mathcal{H}_{k}$. The data sets are either generated with the same mean function and Gaussian noise distribution (for the type $\mathrm{I}$ error), or with a different mean function but the same Gaussian noise distribution (for the type $\mathrm{II}$ error). We bootstrap confidence thresholds and run the test. Figure 6 shows the results for varying values of the data set size $N$. We observe that low values of $N$ lead to an overly aggressive outcome, whereas sufficiently high values yield a conservative threshold as $\alpha$ increases (since the type $\mathrm{I}$ error remains strictly below $\alpha$). However, bootstrapping leads to a sharp decrease in the type $\mathrm{II}$ error, indicating an increased power.
To confirm this increase in power compared to the analytical thresolds, we ran the test with a fixed number of data points but varying measurement noise variances with Gaussian measurement noise, and compared the analytical and bootstrapped thresholds. Importantly, since the output is one-dimensional, the subgaussian noise operator reduces to a scalar, which we set to the ground truth noise variance. Figure 7 shows the results. We see that the analytical thresholds are extremely conservative: they indeed uphold the level, but the errors are effectively independent of $\alpha$. In contrast, bootstrapping preserves the level but is as aggressive as allowed, resulting in a decreasing type $\mathrm{II}$ error as $\alpha$ increases. Overall, this confirms the interest and effectiveness of our bootstrapping scheme.
@@
Figure 7: Type $\mathrm{I}$ and $\mathrm{II}$ errors for the tests with the bootstrapped and analytical thresholds. Analytical thresholds yield a very conservative test (level is less than $\alpha$, but the type $\mathrm{I}$ and $\mathrm{II}$ errors are independent of $\alpha$. In contrast, bootstrapping preserves the guarantee and increases power. The results are averaged over 100 runs of the experiment, with the shaded regions reporting the 2.5% and 97.5% quantiles.
Appendix C Connections between Theorem 4.3 and existing bounds

### v2

@@ A.3.1 Numerical study
We illustrate the preceding ideas by showcasing how the choice of kernel $\kappa$ influences what is being tested, and the power of the test for a fixed amount of data. The metrics and vocabulary used in the setup are identical to those introduced in Section 5, which we recommend to read first.
@@
We consider the Gaussian kernel $k$ on $\mathcal{X}=[-1,1]^{2}$, pick a function in the RKHS of $k$, and generate measurements with two different additive noises: one is $\mathcal{N}(0,s^{2})$, and the other is a mixture of $\mathcal{N}(-\mu,s^{2})$ and $\mathcal{N}(\mu,s^{2})$, for different values of $\mu$ but fixed $s^{2}$. This choice guarantees that the noise distributions have both zero means, but differ for higher moments. We embed the measurements in $\mathcal{G}=\mathcal{H}_{\kappa}$, where $\kappa$ is either the inhomogeneous linear or the Gaussian kernel. The corresponding conditional expectations in $\mathcal{G}$ differ for the Gaussian kernel (since it is characteristic [38]), but coincide for the linear kernel. As a result, a test with such data should trigger with $\kappa$ the Gaussian kernel, and not trigger with $\kappa$ the linear one. We run vector-valued KRR with kernel $K=k\cdot\mathrm{id}_{\mathcal{H}_{\kappa}}$, bootstrap test thresholds, and run the test on the locations present in the data set. Figure 4 shows the resulting empirical positive rate for different $\alpha$-levels, confirming that the Gaussian kernel confidently sees a difference between the distributions, whereas the linear kernel only compares means and therefore does not reject more often than by chance.
@@
However, a rich kernel like the Gaussian one can be less efficient at comparing lower moments of distributions. We show this by comparing empirical error rates of the test for inhomogeneous polynomial kernels of different degrees and Gaussian kernels of different bandwidths for $\kappa$. We use the same setup as before, but this time report type $\mathrm{I}$ and $\mathrm{II}$ errors instead of just the positive rate. For the type $\mathrm{I}$ error, we use the same function from $\mathcal{X}$ to $\mathcal{Z}$ and noise distribution to generate both data sets. For the type $\mathrm{II}$ error, we use different functions from $\mathcal{X}$ to $\mathcal{Z}$, but the same additive Gaussian noise distribution, so that the distributions on $\mathcal{Z}$ differ by their means but higher moments coincide. Figure 5 shows that rich kernels (like the Gaussian one with small bandwidth) — which see fine-grained distributional differences — are less efficient at recognizing differences in the mean than less rich kernels (e.g., linear kernel) tailored to that particular moment of the distribution.
@@
Figure 4: Comparing the Gaussian and inhomogeneous linear kernels on different distributions having the same conditional mean on the measurement set $\mathcal{Z}$. A trigger is a true positive for the Gaussian kernel, and a false positive for the linear one. The results are averaged over 100 runs of the experiment, with the shaded regions reporting the 2.5% and 97.5% quantiles.
@@
Figure 5: Empirical type $\mathrm{I}$ and $\mathrm{II}$ errors for different output kernels. For the type $\mathrm{II}$ error, conditional expectations in $\mathcal{Z}$ differ but higher moments coincide. We see that Gaussian kernels are more conservative and achieve higher type $\mathrm{II}$ error. The results are averaged over 100 runs of the experiment, with the shaded regions reporting the 2.5% and 97.5% quantiles.
@@ Appendix B Bootstrapping
These considerations motivate bootstrapping values of $\beta_{1}$ and $\beta_{2}$ suitable for the problem at hand. Let $D_{1}$ and $D_{2}$ denote the observed data sets of transition pairs from $p_{1}$ and $p_{2}$, with respective sample sizes $n_{1}=\mathrm{len}(D_{1})$ and $n_{2}=\mathrm{len}(D_{2})$. We assume a sufficiently dense finite subset $\mathcal{M}=\{x_{j}\mid j\in\left[J\right]\}\subset\mathcal{S}$, obtained for instance by gridding or random sampling, on which we want the type $\mathrm{I}$ error calibration to hold. We neglect bootstrapping over time, meaning that the resulting values are only meaningful for $(\mathcal{S},n_{1},n_{2})$-guarantees. We further fix $B_{i}=\beta_{i}\cdot\sigma_{D_{i},\lambda_{i}}$ so that the test $\mathcal{T}$ is parameterized by the vector $\beta=(\beta_{1},\beta_{2})\in\mathbb{R}_{>0}^{2}$. Our goal is finding numerically a minimally conservative value for $\beta$ as measured by the empirical type $\mathrm{II}$ error on test scenarios while preserving a type $\mathrm{I}$ error of at most $\alpha$.
In the following, we detail two possible approaches that calibrate $\beta_{1}$ and $\beta_{2}$ independently of each other. The first one essentially bootstraps each value by applying the test on resamplings of the same data set. The second one bootstraps the confidence bounds of the underlying KRR. Both methods work by generating $M$ bootstrap replicates, obtaining statistics $(\bar{\beta}_{i,m})_{m\in\left[M\right]}$ for each $i\in\{1,2\}$, and taking
$\beta_{1}=\text{$(1-\alpha t)$-quantile of }\{\bar{\beta}_{1,m}\},\qquad\beta_{2}=\text{$(1-\alpha(1-t))$-quantile of }\{\bar{\beta}_{2,m}\},$
with a user-specified $t\in(0,1)$. When needed, grid-search over $t$ is inexpensive. We use $t=\frac{1}{2}$ in all simulations.
The following two schemes now differ only in how the families $(\bar{\beta}_{i,m})_{m\in\left[M\right]}$ are generated. Table 1 summarizes the difference in computational complexity.
Naive resampling
@@
A straightforward approach is to resample the data directly. For each $i\in\{1,2\}$, we draw with replacement $M$ pairs of data sets $((\Delta_{m},\Delta_{m}^{\prime}))_{m\in\left[M\right]}$ from $D_{i}$, where each data set in the pair has the same size as $D_{i}$. For each $m$, we then compute $\bar{\beta}_{i,m}$ as the minimal $\bar{\beta}_{m}\in\mathbb{R}_{>0}$ such that
@@
$\left\lVert f_{\Delta_{m},\lambda_{i}}(x)-f_{\Delta_{m}^{\prime},\lambda_{i}}(x)\right\rVert\leq\bar{\beta}_{m}\,\big(\sigma_{\Delta_{m},\lambda_{i}}(x)+\sigma_{\Delta_{m}^{\prime},\lambda_{i}}(x)\big)$
for all $x\in\mathcal{M}$. See Algorithm 2 for more details.
Wild bootstrap
A more scalable alternative is the wild bootstrap, which replaces repeated resampling and re-fitting by random perturbations of the residuals of the nominal KRR estimate. Following Singh and Vijaykumar [27], we adopt anti-symmetric multipliers, which in the scalar-output case were shown to mitigate regularization bias and improve calibration accuracy. Such an analysis is unavailable in our more general setting, but we apply the same construction by straightforward extension to multi-dimensional outputs. However, we use the standard error term $\sigma_{D,\lambda}(x)$ defined in (9) for studentization to remain consistent with the theory-backed form of the confidence bounds identified in Theorem 4.3. For each $i\in\{1,2\}$, draw independent mean-zero multipliers $(q^{(m)}\in\mathbb{R}^{n_{i}})_{m\in\left[M\right]}$ with covariance $I_{n_{i}}-11^{\top}/n_{i}$ and collect the perturbed residual data set
$\Delta_{m}=\left(\left(X_{j},q_{j}^{(m)}\cdot\left(Y_{j}-f_{D_{i},\lambda_{i}}(x_{j})\right)\right)\right)_{(X_{j},Y_{j})\in D_{i}},$
from which $\bar{\beta}_{i,m}$ is obtained as the minimal $\bar{\beta}_{m}\in\mathbb{R}_{>0}$ satisfying
$\left\lVert f_{\Delta_{m},\lambda_{i}}(x)\right\rVert\leq\bar{\beta}_{m}\,\sigma_{\Delta_{m},\lambda_{i}}(x)$
for all $x\in\mathcal{M}$. This method requires only a single KRR fit since the Gram matrix for any $\Delta_{m}$ is the same as for $D_{i}$. See Algorithm 3 for more details.
Table 1: Asymptotic time complexity for computing the test threshold with different calibration methods, in excess of fitting the models $f_{D_{i},\lambda_{i}}$, for $i\in\{1,2\}$, and evaluating the test statistic $\|f_{D_{1},\lambda_{1}}(x)-f_{D_{2},\lambda_{2}}(x)\|$ over all $x\in\mathcal{S}$. Here $n_{i}$ denotes the data set size, $m=|\mathcal{M}|$ the grid size, and $M$ the number of bootstrap replicates.
Method Analytical bound Naive bootstrap Wild bootstrap
Runtime $\mathcal{O}\big(1\big)$ $\mathcal{O}\big(Mn_{i}^{3}+Mmn_{i}^{2}\big)$ $\mathcal{O}\big(Mmn_{i}^{2}\big)$
@@
(a) Influence of data set size $n=n_{1}=n_{2}$.
(b) Influence of regularization parameter $\lambda=\lambda_{1}=\lambda_{2}$.
Figure 6: Type $\mathrm{I}$ and $\mathrm{II}$ errors for the test with different bootstrapping schemes, at varying data set sizes $n$ and regularization levels $\lambda$. Type $\mathrm{I}$ error becomes faithful to the level $\alpha$ when $n$ and $\lambda$ are sufficiently large. Power increases for larger $n$ and smaller $\lambda$. The results are averaged over 100 runs of each experiment. Confidence intervals are omitted for readability. “Singh et al.” refers to the method proposed in [27].
@@
Figure 7: Type $\mathrm{I}$ and $\mathrm{II}$ errors for the test with the bootstrapped and analytical thresholds. Analytical thresholds yield a very conservative test (level is less than $\alpha$, but the type $\mathrm{I}$ and $\mathrm{II}$ errors are independent of $\alpha$. In contrast, bootstrapping preserves the guarantee and increases power. The results are averaged over 100 runs of the experiment, with the shaded regions reporting the 2.5% and 97.5% quantiles.
We assess calibration and power for three bootstrapping schemes: (i) our naive resampling bootstrap; (ii) our wild bootstrap; and (iii) the wild bootstrap of Singh and Vijaykumar [27], which differs from our wild bootstrap only in the studentization term, using
$\sigma_{D_{i},\lambda_{i}}=\left\|\left(\left(k(X_{:n_{i}}^{(i)},X_{:n_{i}}^{(i)})+\lambda_{i}I_{n_{i}}\right)^{-1}k(X_{:n_{i}}^{(i)},x))\right)\mathrm{diag}\left(\varepsilon_{1},...,\varepsilon_{n_{i}}\right)\right\|_{\mathcal{G}^{n_{i}}},$
where $D_{i}=((X_{j}^{(i)},Y_{j}^{(i)}))_{j=1}^{n_{i}}$ is the data set and $\varepsilon_{j}=Y_{j}^{(i)}-f_{D_{i},\lambda_{i}}(X_{j}^{(i)})$ are the residuals as in our wild bootstrap.
We compare the performance of these methods in varying conditions on synthetic data sets generated in a way that either satisfies the null (for the type $\mathrm{I}$ error), or violates it (for the type $\mathrm{II}$ error). For that, we set $\mathcal{X}=[-1,1]^{2}$, take $k$ as the Gaussian kernel with bandwidth $\gamma^{2}=0.25$, and generate data as measurements of functions in $\mathcal{H}_{k}$ with additive Gaussian noise. For the type $\mathrm{I}$ error, we use equal mean functions with Gaussian noise, and for the type $\mathrm{II}$ error use different mean functions with the same noise distribution. For each configuration, we bootstrap the confidence thresholds and run the test.
Varying data set size
Figure 6(a) reports results for data set sizes $n=n_{1}=n_{2}\in\{20,50,100\}$. Type $\mathrm{I}$ error increases with $\alpha$ and improves with $n$. They all uphold the level, except for the naive method with $n=20$. For fixed regularization $\lambda=0.5$, all methods become more conservative as $n$ grows. Accordingly, type $\mathrm{II}$ error improves for increasing $\alpha$ and $n$.
Varying regularization
In Figure 6(b), we now vary the regularization $\lambda=\lambda_{1}=\lambda_{2}\in\{10^{-3},10^{-1},1\}$. At a fixed data set size $n=500$, larger $\lambda$ yields more conservative calibration, while smaller $\lambda$ result in overly aggressive bootstraps and the type $\mathrm{I}$ error exceeding the prescribed level $\alpha$.
Put together, these results indicate that sufficient data and a suitable choice of $\lambda$ are critical for good calibration. Across both experiments, naive resampling is more aggressive (higher type $\mathrm{I}$ and lower type $\mathrm{II}$ error), our wild bootstrap is more conservative (lower type $\mathrm{I}$ and higher type $\mathrm{II}$ error), and the variant by Singh and Vijaykumar [27] lies in between.
Power against analytical calibration
To confirm the increase in power gained by bootstrapping relative to analytical thresholds, we now fix $n$ and $\lambda$, and vary the noise variance $\sigma$. Because the outputs of functions in $\mathcal{H}_{k}$ are one-dimensional, the sub-Gaussian noise operator reduces to a scalar, which we set to the ground-truth variance. For the bootstrapping scheme, we only compare against our naive resampling method. Figure 7 shows that the analytical thresholds are markedly conservative: type $\mathrm{I}$ error remains well below $\alpha$ and both errors are essentially independent of $\alpha$. In contrast, the  type $\mathrm{I}$ error of the bootstrap tracks the level $\alpha$ more faithfully and achieves significantly lower type $\mathrm{II}$ error. Overall, this indicates that bootstrap calibration yields strictly higher power at fixed $\alpha$ without sacrificing control of the type $\mathrm{I}$ error.
For more details on the experiments in this section, see Appendix G.
Appendix C Summary of algorithms
Algorithm 1 summarizes our implementation of the test of Theorem 4.4. The bootstrapping of the multiplicative constants $\beta_{i}$, $i\in\{1,2\}$ it relies on is detailed in Appendix B, and summarized in Algorithms Algorithms˜2 and 3. We emphasize that, in principle, any other bootstrapping algorithm for KRR can be used. All routines assume the form $K=k\cdot\mathrm{id}_{\kappa}$ (c.f Appendix A). Algorithm 4 shows how to compute the test statistic, called the conditional maximum mean discrepancy (CMMD).
Algorithm 1 Kernel conditional two-sample test
1:data sets $D_{1},D_{2}$; tested covariate region $\mathcal{S}\subseteq\mathcal{X}$; bootstrap covariate region $\mathcal{M}\subseteq\mathcal{X}$
[…]
Algorithm 2 Naive bootstrap
1:data set $D$; grid $\mathcal{M}$
2:Chooses $\tilde{\beta}$ s.t. test does not reject on $\mathcal{M}$ for an $(1-\alpha)$-fraction of bootstrapped data sets
3:for $j\leftarrow 1$ to $M$ do
4:  $D_{1},D_{2}\sim D$ $\triangleright$ resample $D_{1},D_{2}$ from $D$ uniformly with replacement
5:  $\mathcal{B}_{j}\leftarrow\max_{x\in\mathcal{M}}\left(\frac{\mathrm{CMMD}(D_{1},D_{2},k,\kappa,\lambda,x)}{\sigma_{D_{1},\lambda}(s)+\sigma_{D_{2},\lambda}(s)}\right)$ $\triangleright$ smallest $\tilde{\beta}$ s.t. test does not reject on $D_{1},D_{2}$
6:return $(1-\alpha)$-quantile of $\{\mathcal{B}_{j}\mid j\in\left[M\right]\}$
Algorithm 3 Wild bootstrap
1:data set $D$; grid $\mathcal{M}$
2:Chooses $\tilde{\beta}$ s.t. test does not reject on $\mathcal{M}$ for an $(1-\alpha)$-fraction of bootstrapped data sets
[…]
Algorithm 4 CMMD: Conditional maximum mean discrepancy
1:covariate $x\in\mathcal{X}$; data sets $D_{1};D_{2}$
[…]
Appendix D Connections between Theorem 4.3 and existing bounds